In one lottery, a player wins the jackpot by matching all five numbers matched by the white balls (1-42) and matching the number on the gold ball (1-35)What is the probability of winning the jackpot.
Answer by fractalier(6550)

The probability can be found by choosing 42 five at a time times 1/35...we get
42!/(42-5)!(5!) * 35 = 850668 * 35 = 29,773,380
That is, one chance in 29,773,380.